OpenStudy (anonymous):

A = 41°, B =61°, c = 19 solve using the law of sines. round to the nearest tenth. please and thank you!

OpenStudy (johnweldon1993):

I thought we already solved this?

OpenStudy (e.cociuba):

Use the Law of Sines.

OpenStudy (anonymous):

@johnweldon1993 nooo, im still lost and never got the solution :/

OpenStudy (e.cociuba):

Hmm maybe @agent0smith can help. He's a good math teacher :)

OpenStudy (agent0smith):

 \[\large \frac{ \sin A }{ a } = \frac{ \sin B }{ b } = \frac{ \sin C }{ c } \]

OpenStudy (anonymous):

hello @agent0smith so i know that a/sin41 b/sin61 and c/sin19 and then sum of a triangles angles is 180 so angle 3 is 78. but where do i go from there?

OpenStudy (agent0smith):

Use the law of sines...  plug in some numbers, pick a side to start with... i'll use C, c and A, a \[\large \frac{19 }{\sin 78 } = \frac{ a }{ \sin 41 }\]